OpenStudy (anonymous):

Another tough one. Not sure how to work this. I know the formula for it will be expressed as C= K*P1*P2/d^2 The average number of daily phone calls, c, between two cities varies jointly as the product of their populations P1 and P2 and inversely as the square of the distance, d, between them. Given the distance between City A (population 777,000) and City B(population 3,695,000) is 420 and that the average number of daily phone calls between them is 427,000 what is the value of k?

OpenStudy (anonymous):

not necessarily we could all be average

OpenStudy (anonymous):

C = 427,000 P1 = 777,000 P2 = 3,695,000 d = 420 Plug and chug

OpenStudy (anonymous):

\[C=\frac{ kP _{1}P _{2} }{ d ^{^{2}} }\] Is the formula more exactly

OpenStudy (anonymous):

\[427,000=\frac{k\times 777,000\times 3,695,000}{420^2}\]

OpenStudy (anonymous):

So does k come to equal out to be when rounded by two decimal places to .03 from 0.0262355996 ?

OpenStudy (anonymous):

0.03 is correct if rounding to two decimal places.

OpenStudy (anonymous):

Based on the longer number you gave below it

OpenStudy (anonymous):

Okay but is the longer number right?

OpenStudy (anonymous):

I don't have a calculator on me, but if you solved for K by multiplying and dividing correctly, that should be it.

OpenStudy (anonymous):

Found a calculator, your answer is correct!

OpenStudy (anonymous):

Thank you! Appreciate the help. :-)

OpenStudy (anonymous):

No problem!